NYSDEC Oversight

The NYSDEC provides oversight of activities conducted by the major responsible parties at various Greenpoint petroleum remediation sites in Brooklyn, New York. NYSDEC’s subcontractor (Ecology and Environment Engineering P.C.) conducts bi-weekly field oversight of PRP activities to assist NYSDEC in developing a better understanding of the day to day investigation and remedial activities that occur in the Greenpoint area and report any pertinent issues to NYSDEC’s project manager.

The following sections summarize recent activities observed at the various Greenpoint remediation sites in Brooklyn, New York.


Date of Oversight Observation: June 13, 2018

Former ExxonMobil Terminal

Description of Activities

  1. Roux informed EEEPC that the first round of on-site DMR  sampling was performed on 6/5/18.
  2. Roux informed EEEPC that the second quarter gauging event was conducted on 5/30/18.
  3. Roux informed EEEPC that a subcontractor will be on site next week to clean out the 20,000 gallon frac tank adjacent to the treatment building.
  4. Roux informed EEEPC that starting on 7/16/18 and to coincide with a 6 to 8-week shutdown of the off-site free-product recovery system (ORS) facility, ORS recovery wells recovering higher amounts of product will be reprogrammed to flow to the free-product recovery and containment system (RCS) facility while the ORS is offline. Roux also in-formed EEEPC that this work will require some software changes and turning on valves along the pipelines that are already in place, so no new ground intrusive work is required.
  5. EEEPC observed the on-site treatment system in operation and recorded current operating parameters of the air strippers from the control screen.
  6. EEEPC and Roux collected a sample of the on-site effluent for visual obsevation and compared it to a previously collected sample taken from Newtown Creek in the vicinity of the outfall.

 Planned Activities

  1. Routine treatment system maintenance.

ExxonMobil Off-Site Plume Area

Description of Activities

  1. Roux informed EEEPC that the first round of off-site DMR sampling was performed on 6/12/18.
  2. Roux informed EEEPC that the second round of off-site DMR sampling is scheduled for 6/17/18.
  3. Roux informed EEEPC that the ORS treatment system will be shut down on 7/16/18 for 6 to 8 weeks. During the shutdown, piping and pumps will be reconfigured in order to improve flow rate and flow efficiency through the system. Additionally, portions of pipe will be replaced with larger diameter pipe, and the process tank and pipes will be cleaned.
  4. Roux informed EEEPC that the treatment facility is offline today to change out a valve in the sand filter line and change out belts in the blower unit. The facility will be back online later today.

Planned Activities

  1. Routine treatment system maintenance.

KM Terminal (Former BP Terminal)

Description of Activities

  1. Antea informed EEEPC that treatment system bag filters were changed out on 6/11/18. Antea also in-formed EEEPC that the 50 and 100 micron bag filters they are now using last longer and require less frequent change-outs than previously used filters.
  2. Antea informed EEEPC that starting on 5/11/18, the treatment system is now running on 50 and 100 micron bag filters. These sizes were tested on the system by Antea and met both the removal and flow criteria.
  3. Antea informed EEEPC that since last Friday (6/8/18), all recovery wells are now operating on compressed air; nitrogen tanks are no longer used or will be delivered to the site.
  4. EEEPC and Antea collected a sample of on-site effluent for visual observation and compared it to a previously collected sample taken from Newtown Creek.

 Planned Activities

  1. None.

Former Paragon Oil Terminal/Apollo Street Creek Parcels

Description of Activities

  1. Leidos informed EEEPC that a subcontractor will be completing the elevated walkway for the tallest tank on the middle bulkhead by this week and work will continue into next week. 
  2. Leidos informed EEEPC that construction for the load plates to support the new housing structure will begin next week. 
  3. Leidos informed EEEPC the barge float-in of the new housing structure is tentatively scheduled for 7/11/18, with 7/12/18 as a back-up date. 
  4. Leidos informed EEEPC that a subcontractor performed a creek-side bulkhead inspect last week and repaired six (6) seams as part of preventative work. 
  5. Leidos informed EEEPC that the new sub-slab depressurization system (SSDS) for the office space on 50 Bridgewater Street has been online and working since 5/28/18. 
  6. Leidos informed EEEPC that Empire Merchants will have a 2-week shutdown starting on 7/16/18. 

 Planned Activities

  1. Routine treatment system maintenance.
  2. Continuation of elevated walkway construction.

Newtown Creek Observations

Former Paragon Terminal Bulkhead
EEEPC observed conditions along the upper, middle, and lower bulkheads, within the containment areas, and in Newtown Creek in the vicinity of the bulkhead. A 1-foot long petroleum sheen was observed inside the lower bulkhead boom containment area in Newtown Creek. Sediment plumes were observed inside of the middle and lower bulkhead boom containment areas as well.

Meeker Avenue Bulkhead
EEEPC observed the conditions of Newtown Creek at the end of Meeker Avenue in the vicinity of the ExxonMobil Off-site outfall and adjacent storm outfalls. Storm outfall D was was discharging a small amount of water. No sheens or sediment plumes were observed in the vicinity or within visible distance upstream or downstream of the outfalls. This observation took place approximately 1 hour and 20 minutes before the 9:48 AM 4.4’ AMSL high tide.


This page was last updated on Wednesday, June 20, 2018.